Using sum of the years-digits-method compute the depreciation of equipment during the third year of its eight years estimated life. The brandnew cost of the equipment is 20,000 and has a salvage value of 2,500
Answer #1

Depreciable amount = Cost-Salvage Value = 20000-2500 = 17500

Life is 8 years so sum of years = 1+2+3+4+5+6+7+8 = 36

During the third year remaining life of the asset is 8-2 = 6 years

Depreciation = Remaining life/ sum of years * Depreciable amount = 6/36 * 17500 = 17500/6 = 2916.67

Hence 2916.67 is answer